1. 3n=(6-n)(-3)
2. -10= 4- 7x/4
3. 1/4(y+8)=5
4 7.6x + 3.7= 2.8-1.6x

Answer by checkley75(3666)

1) 3N=(6-N)(-3), 3N=-18+3N, 3N-3N=-18, 0N=18 NO SOLUTION FOR N
2) -10=4-7X/4, -10+4=-7X/4, 6=-7X/4, CROSS MULTIPLY: -7X=24, X=24/-7, X=-24/7
3) 1/4(Y+8)=5, Y/4+8/4=5, Y/4=5-8/4, Y/4=5-2, Y/4=3, Y=12
4) 7.6X+3.7=2.8-1.6X, 7.6X+1.6X=2.8-3.7, 9.2X=-.9, X=-.9/9.2